OpenStudy (anonymous):

divide 6a^3b^2/9a^5b^4/3a^2b^2

OpenStudy (anonymous):

@MathLegend for my final i got 2a/3a^3b^4 is this right?

OpenStudy (mathlegend):

Is this a fraction?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

sorry it took me long i had to step away fromth e computer for a minute

OpenStudy (mertsj):

\[\frac{6a^3b^2}{\frac{9a^5b^4}{3a^2b^2}}\]

OpenStudy (mertsj):

Is that the problem?

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

Then simplify the denominator first: \[\frac{6a^3b^2}{3a^3b^2}=2\]

OpenStudy (anonymous):

2

OpenStudy (mertsj):

Yep

OpenStudy (anonymous):

wait it's supposed to be 3a^2b^2 not 3a^3b^2

OpenStudy (anonymous):

so i have 2a/3a

OpenStudy (anonymous):

is this correct

OpenStudy (mertsj):

no

OpenStudy (mertsj):